OpenStudy (weemzman):

What is the value of b2 - 4ac for the following equation? 2x^2 + 3x = -1

sam (.sam.):

\[2x^2 + 3x = -1\] \[2x^2+3x+1=0 ~~\leftarrow ~Relate~this~\to~~ax^2+bx+c=0\]

sam (.sam.):

a=2 b=3 c=1

sam (.sam.):

\[b^2-4ac\] Put the values in and you're done